# The impact of data set similarity and diversity on transfer learning success in time series forecasting

### v1

@@ 3.1 Source data
As source data for pre-training, we select five data sets from the Monash Time Series Forecasting Repository (Godahewa et al.,, 2021): M5 (Makridakis et al.,, 2022), M4 (Spiliotis et al.,, 2020), Electricity (Lai et al.,, 2018), NN5 (Taieb et al.,, 2012) and Exchange Rate (Lai et al.,, 2018). In that, we follow the selection criteria public availability, domain and size diversity, and the availability for download within the gluonts framework. We directly use the available weekly NN5 and M4 data sets and otherwise aggregate the time series to weekly granularity. To keep the scope of the experiment manageable, we disregard other granularities and restrict to univariate time series. The chosen source data sets and their size are listed in Table 1.
For each time series, we use 80% of the time steps as training set to pre-train the source models. Overall, we consider around 31,300 time series with 7,5 million time steps for pre-training.
Data set
Domain
Description
Time Series
Training Time Steps
Training Size
M5
Sales
Walmart SKU Sales
30,490
219
6.7 mio
M4
@@
diverse series
359
2,000
718,000
Electricity
Energy
electricity consumption
321
102
32,700
NN5
Banking
cash withdrawals at ATMs
111
90
10,000
Exchange Rate
Banking
exchange rates
8
996
8,000
Sum
31,300
3,400
7.5 mio
@@ 3.2 Target data
To test the transferability of knowledge stored in pre-trained models, we select three publicly available target data sets based on the criteria applied to select source data sets: Kaggle Web Traffic (Godahewa et al.,, 2020), Traffic (Lai et al.,, 2018) and Solar Energy (Lai et al.,, 2018). In addition, we consider two SKU wholesales project data sets from two different wholesalers, Wholesaler1 and Wholesaler2, a domain without well-known public data. Again, we select the data sets aggregated to weekly granularity, if available or aggregate by ourselves. Table 2 features a short description per target data set. Note that the target domains differ from those of the source data to examine transfer learning, data set similarity and diversity more broadly. The energy domain appears in the sources and in the targets, but the data generating processes differ.
Again, we split the data sets 80:20 along the time axis to train the scratch model or fine-tune the pre-trained model, and test the model. To keep the computational effort of the experiment manageable, we restrict the target test data sets for evaluation to a maximum of 1,000 randomly drawn time series each. For data sets containing fewer time series, we account for the entire data set. Depending on the size of the test data set, a different number of forecasts is evaluated for the targets, gained in a rolling origin manner (Tashman,, 2000). For the Solar Energy data set, the test set is larger than 20%, since we need at least one forecast for evaluation.
Data set
Domain
Description
Time Series
Training Time Steps
Training Size
Test Time Steps
Test Size
Rolling Origins
Wholesaler1
Sales
SKU sales
113,000
130
15 mio
32
32,000
18
Wholesaler2
Sales
SKU sales
55,000
124
7 mio
31
31,000
17
Kaggle Web Traffic
Web
Wikipedia web pages
145,000
91
13 mio
23
23,000
9
Taffic
Transport
Occupancy rate of care lanes
862
68
59,000
17
16,000
3
Solar Energy
Energy
Solar power production of a state
137
35
5,000
15
2,000
1
Sum
314,000
448
35 mio
118
104,000
48
@@ 3.3 Transfer learning and scratch models
We apply the neural network architecture DeepAR (Salinas et al.,, 2020) as the source and scratch model architecture. DeepAR is well known in the forecasting community, can be trained in a global manner, i.e. over multiple time series to extract joint knowledge, and is able to output probabilistic forecasts, i.e. gives an estimation about the forecast uncertainty.
For robustness, we use an ensemble composed of five DeepAR models. Pre-trained and scratch models share the hyper-parameters, listed in Table 3. The models encompass as many output nodes as the forecast horizon for direct forecasting. Given the implicit generality of transfer learning, we waive hyper-parameter optimization and take the default values of the implementations in the gluonts (Alexandrov et al.,, 2020) package. One exception is the distribution of the DeepAR, where we select the negative binomial as proposed in (Salinas et al.,, 2020). The gluonts DeepAR model comes with a default input scaling to the batch mean which helps in adapting to the target data sets’ scales. Hence, we do not scale the data during pre-processing.
Scratch training, pre-training and fine-tuning involves all layers and is conducted with the gluonts mx.trainer with an early stopping adaptation, including a maximum of 800 epochs.
@@
3.4 Statistical models
For the comparison of the forecasting performance with statistical benchmarks, we implement the ETS model via the gluonts R wrapper, r_forecast, which uses the R forecast package (Hyndman et al.,, 2024). There, the default parameters are set, which corresponds to an automatically selected error, trend and season type, an estimation of the model parameters and trying both, damped and non-damped trend. The model can also output prediction intervals by sampling from the error distribution. The performance of the Naive forecast is included indirectly and only during training time steps in the following metrics to get relative ones, but not explicitly described in the result analysis.
@@ 3.5 Metrics
For the evaluation of forecast performance, we take the accuracy, bias, uncertainty estimation and training time into account. For any target data set, we denote the number of time series with $m$, and the series themselves with $y^{i},i=1,...,m$. The test set contains rolling origins $r=1,...,R$, which start after the past time steps $t=1,...,T_{r}$ and encompass the test time steps $t=T_{r}+1,...,T_{r}+h$ during the forecast horizon $h$. The accuracy and bias measure refers to the median forecast $\hat{y}^{i}_{t}$ of a time series $i$ at time $t$. The uncertainty estimation error refers to the forecast quantiles $\hat{q}^{i}_{j,t}$, where $j$ is the quantile percentage of time series $i$ at time $t$.
For the accuracy we consider the Average Relative Root Mean Squared Error (AvgRelRMSE) given in Eq. 1 in analogy to the relative measures in Hewamalage et al., (2023). It is the average over rolling origins and time series of the forecast RMSE in Eq. 2 relative to the naive forecast RMSE during past time steps in Eq. 3. Note, that we take the RMSE instead of the MAE for consistency to measure the error of the naive forecast.
$AvgRelRMSE$ $=\frac{1}{R}\sum_{r=1}^{R}\left(\frac{1}{m}\sum_{i=1}^{m}(RMSE^{i,r}/RMSE_{naivepast}^{i,r})\right)$ (1)
$RMSE^{i,r}$ $=\sqrt{\frac{1}{h}\sum_{t=T_{r}+1}^{T_{r}+h}(y_{t}^{i}-\hat{y}_{t}^{i})^{2}}$ (2)
$RMSE_{naivepast}^{i,r}$ $=\sqrt{\frac{1}{T_{r}-1}\sum_{t=2}^{T_{r}}(y_{t}^{i}-y_{t-1}^{i})^{2}}$ (3)
@@
$ME$ $=\frac{1}{R}\sum_{r=1}^{R}\left(\frac{1}{m}\sum_{i=1}^{m}(ME^{i,r})\right)$ (4)
$ME^{i,r}$ $=\frac{1}{h}\sum_{t=T_{r}+1}^{T_{r}+h}(y_{t}^{i}-\hat{y}_{t}^{i})$ (5)
We assess the uncertainty estimation of the sample forecasts by a modified Mean Scaled Interval Score (MSIS) (Gneiting and Raftery,, 2007) with $\alpha=0.05$, given in Eq. 6. It averages the MSIS of each forecast in Eq. 7, which again uses the RMSE of the naive forecast in the past instead of its MAE in the denominator.
$MSIS$ $=\frac{1}{R}\sum_{r=1}^{R}\left(\frac{1}{m}\sum_{i=1}^{m}(MSIS^{i,r})\right)$ (6)
$MSIS^{i,r}$ $=\frac{\frac{1}{h}\sum_{t=T_{r}+1}^{T_{r}+h}(\hat{q}^{i}_{97.5,t}-\hat{q}^{i}_{2.5,t})+\frac{2}{0.05}(\hat{q}^{i}_{2.5,t}-y_{t}^{i})\mathds{1}_{(y_{t}^{i}<\hat{q}^{i}_{2.5})}+\frac{2}{0.05}(y_{t}^{i}-\hat{q}^{i}_{97.5,t})\mathds{1}_{(y_{t}^{i}>\hat{q}^{i}_{97.5,t})}}{RMSE_{naivepast}^{i,r}}$ (7)
To compare training times on the target data, we report the training times of the scratch model and of fine-tuning the pre-trained models on a Linux server with 32 CPUs and 77 GB RAM. For that, we integrated the python package codecarbon (Schmidt et al.,, 2021).
@@ 4 Result analysis
In the following, the forecasting performance of the pre-trained, scratch and benchmark models within the experimental framework described in Section 3 is evaluated regarding accuracy, bias, uncertainty estimation and training time with the help of the defined metrics in Subsection 3.5.
Source / Target Solar Traffic Kaggle Wholesaler2 Wholesaler1
@@
Table 4: Zero-shot/fine-tuned transfer learning, scratch and benchmark accuracy performance in AvgRelRMSE (see Eq. 1). Smaller values are better. The best results for each target data set are printed in bold and best results within the pre-trained models are underlined.
@@ 4.1 Accuracy
Table 4 shows the accuracy measured in terms of the Average Relative RMSE (see Eq. 1) of the pre-trained, scratch and benchmark models for the different source and target combinations. There is always at least one pre-trained model that beats the scratch model in the zero-shot case, i.e. without having ever seen the target data and which has been trained on data of another domain. The ETS model is the best fit for Traffic and Wholesaler1 and cannot be beaten even after fine-tuning.
Within the pre-trained models, the Multisource and M4 source models achieve the best transfer learning accuracy. Fine-tuning enhances the accuracy of the pre-trained models in 74% of the cases. This particularly applies to the Electricity source data set, where accuracy improves from the worst in zero-shot forecasting to intermediate after fine-tuning. In contrast, the Multisource and the M4 source models tend to get worse with fine-tuning. While the M5 data is a huge part of Multisource (89%), their respective source models provide distinct accuracy measures on the targets. The Multisource model displays a superior performance than the Ensemble, although both were trained on the same data. In contrast to the other target data sets, pre-training on the M5 source provides better zero-shot performance on the wholesales data sets, which might be due to similar time series characteristics in retail and wholesales. The last observation is also in line with the literature in Section 2.
Figure 1: Zero-shot/scratch (red circles) and fine-tuned (blue circles) accuracy in AvgRelRMSE (see Eq. 1) per source model across the five target data sets.
Figure 1 compares each model’s zero-shot and fine-tuned accuracy given one source data set across the different target data sets. This view demonstrates the overall effect of fine-tuning leading to enhanced accuracy, while the performance on the Traffic data set often constitutes an outlier with higher error. The model pre-trained on the Multisource shows similarly good performance in both cases and on four out of five target data sets.
@@ 4.2 Bias
Source / Target Solar Traffic Kaggle Wholesaler2 Wholesaler1
@@
Table 5 reports the scaled bias in terms of the Mean Error (see Eq. 4) for the zero-shot and fine-tuned case. To compare the zero-shot and fine-tuned bias, we scale the bias over the results of both modes per target data set, by dividing by the standard deviation. For both modes, with one exception, there is always a pre-trained model which reveals a better bias than the Scratch and the ETS model. Overall, fine-tuning reduces the forecasting bias in 51% of the time. Within the pre-trained models, the one pre-trained on the Exchange Rate provides the lowest bias.
To summarize: in general, the pre-trained models provide a better bias than the Scratch and the benchmark model.
@@ 4.3 Uncertainty estimation
Source / Target Solar Traffic Kaggle Wholesaler2 Wholesaler1

### v2

@@ 3.1 Source data
As source data for pre-training, we select five publicly available data sets from the Monash Time Series Forecasting Repository (Godahewa et al., 2021) which are diverse regarding size and domain: M5 (Makridakis et al., 2022), M4 (Makridakis et al., 2020), Electricity (Lai et al., 2018), NN5 (Crone, 2009) and Exchange Rate (Lai et al., 2018). We directly use the available weekly NN5 and M4 data sets and otherwise aggregate the time series to weekly frequency. To keep the scope of the experiment manageable, we disregard other frequencies and restrict to univariate time series. The chosen source data sets and their size are listed in Table 1. For each time series, we use 80% of the time steps as training set to pre-train the source models. Overall, we use 31,300 time series, and in sum around 7.5 million observations for pre-training.
Data set Domain Description # Time Series # Training Periods Training Size
M5 Sales Walmart SKU sales 30,490 219 6,677,310
M4 Diversea Diverse series 359 2,000 718,000
Electricity Energy Electricity consumption 321 102 32,742
NN5 Banking Cash withdrawals at ATMs 111 90 9,990
Exchange Rate Banking Exchange rates 8 996 7,968
Sum 31,289 3,407 7,446,010
  • a
@@ 3.2 Target data
To test the transferability of knowledge stored in pre-trained models, we select three publicly available target data sets based on the criteria applied to select source data sets: Kaggle Wiki  (Anava et al., 2017), Traffic (Lai et al., 2018) and Solar Energy (Lai et al., 2018). In addition, we consider two SKU wholesales project data sets from two different wholesalers, Wholesaler1 and Wholesaler2, a domain without well-known public data. Again, we select the data sets aggregated to weekly frequency, if available or aggregate by ourselves. Table 2 features a short description per target data set. Note that the target domains differ from those of the source data to examine transfer learning, data set similarity and diversity more broadly. The energy domain appears in the sources and in the targets, but the underlying business processes differ.
Again, we split the data sets 80:20 along the time axis to train the Scratch model or fine-tune the pre-trained model, and test the model. To keep the computational effort of the experiment manageable, we restrict the target test data sets for evaluation to a maximum of 1,000 randomly drawn time series each. For data sets containing fewer time series, we account for the entire data set. We evaluate the forecasts on rolling origins (Tashman, 2000) of the target test sets. For the Solar Energy data set, 20% of the time steps would be less than the forecast horizon. Since we need at least one forecast per time series for evaluation, we enlarge the test set here.
Data set Domain Description Time Series Training Periods Training Size Test Periods Test Size Rolling Origins
Wholesaler1 Sales SKU sales 113,000 130 14,690,000 32 32,000 18
Wholesaler2 Sales SKU sales 55,000 124 7,820,000 31 31,000 17
Kaggle Wiki Web Web traffic of Wikipedia pages 145,000 91 13,195,000 23 23,000 9
Traffic Transport Occupancy rate of car lanes 862 68 58,616 17 14,654 3
Solar Energy Energy Solar power production of a state 137 35 4,795 15 2,055 1
Sum 313,999 448 35,768,411 118 102,709 48
@@ 3.3 Transfer learning and scratch models
We apply the neural network architecture DeepAR (Salinas et al., 2020) as the source and Scratch model architecture. DeepAR is well known in the forecasting community and can be trained in a global manner, i.e. over multiple time series to extract joint knowledge. Further, it is able to output probabilistic forecasts, i.e. give an estimation about the forecast uncertainty.
For robustness, we use an ensemble composed of five DeepAR models. Pre-trained and Scratch models share the hyper-parameters, listed in Table 3. The models encompass as many output nodes as the forecast horizon for direct forecasting. To save computational resources, we waive hyper-parameter optimization and take the default values of the implementations in the gluonts package (Alexandrov et al., 2020). One exception is the output distribution of the DeepAR, where we select the negative binomial as proposed in (Salinas et al., 2020). The DeepAR model comes with a default item-wise mean scaling, which helps in adapting to the target data sets’ scales. Hence, we do not scale the data during pre-processing.
Scratch training, pre-training and fine-tuning involve all layers and are conducted with the gluonts mx.trainer with an early stopping adaptation, including a maximum of 800 epochs.
@@
3.4 ETS benchmark model
As representative local statistical benchmark, we use the ETS model via the gluonts R wrapper, r_forecast, which uses the R forecast package (Hyndman et al., 2024). There, the default parameters are set, which corresponds to an automatically selected error, trend and season type, an estimation of the model parameters and trying both, damped and non-damped trend. Depending on the model type, prediction intervals are gained according to closed form forecast variance formulas if available for the respective ETS model or simulated sample paths otherwise  (Hyndman and Athanasopoulos, 2021).
@@ 3.5 Metrics
For the evaluation of forecast performance, we take the accuracy, bias, and uncertainty estimation into account. For any target data set, we denote the number of time series with $m$, and the series themselves with $y_{i},i=1,...,m$. The test set contains rolling origins $r=1,...,R$, which start after the past time steps $t=1,...,T_{r}$ and encompass the test time steps $t=T_{r}+1,...,T_{r}+h$ during the forecast horizon $h$. The accuracy and bias measure refers to the median forecast $\hat{y}_{i,t}$ of a time series $i$ at time $t$. The uncertainty estimation error refers to the forecast quantiles $\hat{q}_{i,j,t}$, where $j$ is the quantile percentage of time series $i$ at time $t$.
For the accuracy we consider the Average Root Mean Squared Scaled Error (AvgRMSSE) given in Eq. 1 which is based on the RMSSE (Makridakis et al., 2022). It is the average over rolling origins and time series of the forecast RMSSE in Eq. 2 relative to the naive forecast RMSE in the training data Eq. 3.
$AvgRMSSE$ $=\frac{1}{R}\sum_{r=1}^{R}\left(\frac{1}{m}\sum_{i=1}^{m}(RMSE_{i}^{r}/RMSE_{i,naivepast}^{r})\right)$ (1)
$RMSE_{i}^{r}$ $=\sqrt{\frac{1}{h}\sum_{t=T_{r}+1}^{T_{r}+h}(y_{i,t}-\hat{y}_{i,t})^{2}}$ (2)
$RMSE_{i,naivepast}^{r}$ $=\sqrt{\frac{1}{T_{r}-1}\sum_{t=2}^{T_{r}}(y_{i,t}-y_{i,t-1})^{2}}$ (3)
@@
$ME$ $=\frac{1}{R}\sum_{r=1}^{R}\left(\frac{1}{m}\sum_{i=1}^{m}(ME_{i}^{r})\right)$ (4)
$ME_{i}^{r}$ $=\frac{1}{h}\sum_{t=T_{r}+1}^{T_{r}+h}(y_{i,t}-\hat{y}_{i,t})$ (5)
We assess the uncertainty estimation of the sample forecasts by a modified Mean Scaled Interval Score (MSIS) (Makridakis et al., 2020), given in Eq. 6. It averages the MSIS with $\alpha=0.05$ of each forecast in Eq. 7 over time series and rolling origins. For consistency in scaling, we use the RMSE of the naive forecast in the past (Eq. 3) instead of its MAE, as proposed by the authors, in the denominator.
$MSIS$ $=\frac{1}{R}\sum_{r=1}^{R}\left(\frac{1}{m}\sum_{i=1}^{m}(MSIS_{i}^{r})\right)$ (6)
$MSIS_{i}^{r}$ $=\frac{\frac{1}{h}\sum_{t=T_{r}+1}^{T_{r}+h}(\hat{q}_{i,97.5,t}-\hat{q}_{i,2.5,t})+\frac{2}{0.05}(\hat{q}_{i,2.5,t}-y_{i,t})\mathds{1}_{(y_{i,t}<\hat{q}_{i,2.5,t})}+\frac{2}{0.05}(y_{i,t}-\hat{q}_{i,97.5,t})\mathds{1}_{(y_{i,t}>\hat{q}_{i,97.5,t})}}{RMSE_{i,naivepast}^{r}}$ (7)
We report fine-tuning times of pre-trained models in relation to scratch training times on a Linux virtual machine with 32 CPUs and 77 GB RAM in the Appendix Table 13. For that, we use the python package codecarbon (Schmidt et al., 2021).
@@ 4 Result analysis
In the following, we provide the accuracy, bias, and uncertainty estimation performance of the pre-trained, Scratch and benchmark model.
Source / Target Solar Traffic Kaggle Wiki Wholesaler2 Wholesaler1
@@
Table 4: Zero-shot/fine-tuned transfer learning, scratch and benchmark accuracy performance in AvgRMSSE (see Eq. 1). Smaller values are better. The best results for each target data set and mode are printed in bold and best results within the pre-trained models are underlined.
@@ 4.1 Accuracy
Table 4 shows the accuracy measured in terms of the Average RMSSE (see Eq. 1) of the pre-trained, Scratch and benchmark models for the different source and target combinations. There is always at least one pre-trained model that beats the Scratch model in the zero-shot case, i.e. without having ever seen the target data and which has been trained on data of another domain. In 40% of the time zero-shot pre-trained models beat the Scratch model. If a pre-trained model provides a zero-shot performance that is superior to the Scratch model, only in 35% it gets even better with fine-tuning. The ETS model is the best fit for Traffic and Wholesaler1 and cannot be beaten even after fine-tuning.
The fine-tuned pre-trained models beat the Scratch model in 89% of the cases, showing the potential of pre-trained models. Note that the shared hyperparameters of pre-trained and Scratch models are default, making their practical usage realistic. This is in contrast to other papers comparing huge architectures used for pre-training also for Scratch model training on a single data set. Within the pre-trained models, the Multisource and M4 source models achieve the best transfer learning accuracy. Fine-tuning enhances the accuracy of the pre-trained models in 74% of the cases. This particularly applies to the Electricity source data set, where accuracy improves from the worst in zero-shot forecasting to intermediate after fine-tuning. In contrast, the Multisource and the M4 source models tend to get worse with fine-tuning. A possible reason might be that the model parameters converge against those of the Scratch model during fine-tuning. If this is worse than the zero-shot performance, the fine-tuned performance drops. While the M5 data is a huge part of Multisource (89%), their respective source models provide distinct accuracy measures on the targets. The Multisource model displays a superior performance than the Ensemble, although both were trained on the same data. In contrast to the other target data sets, pre-training on the M5 source provides better zero-shot performance on the wholesales data sets, which might be due to similar time series characteristics in retail and wholesales. This is also in line with the literature in Section 2.
Figure 2: Zero-shot/scratch (red circles) and fine-tuned (blue circles) accuracy in AvgRMSSE (see Eq. 1) per source model across the five target data sets. View in color.
Figure 2 compares each model’s zero-shot and fine-tuned accuracy given one source data set across the different target data sets. This view demonstrates the overall effect of fine-tuning leading to enhanced accuracy. The model pre-trained on the Multisource shows similarly good performance in both cases and on four out of five target data sets.
@@ 4.2 Bias
Source / Target Solar Traffic Kaggle Wiki Wholesaler2 Wholesaler1
@@
Table 5 reports the scaled bias in terms of the Mean Error (see Eq. 4) for the zero-shot and fine-tuned case. To compare the zero-shot and fine-tuned bias, we scale the bias over the results of both modes per target data set, by dividing by the bias standard deviation. For both modes, with one exception, there is always a pre-trained model which reveals a better bias than the Scratch and the ETS model. Overall, fine-tuning reduces the forecasting bias in 51% of the time. Within the pre-trained models, the one pre-trained on the Exchange Rate provides the lowest bias.
To summarize: depending on the source data set, pre-trained models reduce bias compared to the Scratch and the benchmark model.
@@ 4.3 Uncertainty estimation
Source / Target Solar Traffic Kaggle Wiki Wholesaler2 Wholesaler1